# Convergence Analysis of Randomized Subspace Normalized SGD under Heavy-Tailed Noise

### v1

@@ 1 Introduction
Table 1: Convergence guarantees for stochastic gradient methods with randomized subspace updates (RS) and direction normalization (Norm). Exp./w.h.p.: in expectation/high probability. Noise: sub-Gaussian, BV, and $p$-BCM; see Assumptions 2.5, 2.4 and 2.6, respectively.
@@
However, these light-tailed noise conditions can be overly restrictive in modern machine-learning workloads, where stochastic gradients are often heavy-tailed [27]; this limitation is not caused by the randomized subspace projection but rather reflects the well-known sensitivity of SGD-type updates to heavy-tailed gradient noise [23, 10]. To overcome this issue, we leverage normalized stochastic gradient methods (NSGD), which admit both in-expectation and high-probability convergence guarantees under the weaker bounded $p$-th moment assumption stated later in Assumption 2.6 [11]. Unlike clipping-based methods, whose analyses typically require problem- or noise-dependent parameters, including the tail index $p$ [34, 22], NSGD can be analyzed without knowing $p$.
@@ 2.3 Problem Setting and Assumptions
While Assumption 2.5 is standard, it can be too restrictive in regimes where the stochastic gradient noise is heavy-tailed. To accommodate such behavior, we allow the noise to have only a bounded $p$-th moment for some $p\in(1,2]$.
@@
These noise assumptions are related as follows: [^3] Reason. Assumption 2.4 implies Assumption 2.5 by Jensen’s inequality. Assumption 2.5 is Assumption 2.6 with $p=2$, and Assumption 2.6 for some $p$ implies Assumption 2.6 for any $q\leq p$ (again by Jensen). Hence, Assumption 2.5 implies Assumption 2.6 for every $p\in(1,2]$.
@@ 4 High-Probability Convergence of RS-SGD
Assume that the stochastic gradient $\nabla f(x,\xi)$ satisfies Assumptions 2.3 and 2.4. For a fixed $x\in\mathbb{R}^{d}$, define the mini-batch gradient estimator
@@
By Lemma 4.1, increasing the mini-batch size reduces the noise, and the following theorem establishes convergence.
@@
Suppose that Assumptions 2.1, 2.2, 2.3 and 2.4 hold. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by RS-SGD. Let $\bar{B}\coloneqq\lceil\max\{1,BT^{q}\}\rceil$ for $B>0$ and $q>0$, $\alpha\coloneqq 24\sqrt{\frac{e(e+1)}{2}}$, and $\bar{\eta}=\frac{r}{Ld}$. Further define
@@ 5.1 In-expectation Bound
When bounding the first term in Proposition 5.2 by $\|S\|$, the relationship between $\operatorname{tr}(S)$ and $\|S\|$ becomes important. To this end, we introduce the following quantity.
@@
Suppose that Assumptions 2.1, 2.2, 2.3 and 2.6 hold. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by RS-NSGD. We set
@@ Appendix B Missing Proofs for RS-SGD
Assume $\mathbb{E}[X]=0$ and $\|X\|$ is $\sigma$-sub-Gaussian in the sense of Definition B.1. Then for all $u\in\mathbb{R}^{d}$,
@@
By Definition B.1 with $\lambda=1/\sigma$, $\mathbb{E}[\exp(\|X\|^{2}/\sigma^{2})]\leq e$. Therefore,
@@
Proposition B.4 (Generalization of Lemma 4.1).
@@
By Lemma B.2, for each $i$ and all $u\in\mathbb{R}^{d}$,
@@
Thus, the assumption of Lemma B.3 holds for $\bar{X}$ with
@@
Applying Lemma B.3 yields
@@
Assume $F$ is $L$-smooth and bounded below by $F^{\ast}$. Assume $\mathbb{E}[\nabla f(x,\xi)]=\nabla F(x)$ and that $\|\nabla f(x,\xi)-\nabla F(x)\|$ is $\sigma$-sub-Gaussian (in the sense of Definition B.1). Let $P_{k}$ be i.i.d. Haar random matrices, and let $\{x_{k}\}_{k\geq 0}$ be generated by the update rule
@@
Suppose that $F$ satisfies Assumptions 2.1 and 2.2, and that the gradient satisfies Assumptions 2.3 and 2.4. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-SGD iteration. Let $\bar{B}\coloneqq\lceil\max\{1,BT^{q}\}\rceil$ for $B>0$ and $q>0$, $\alpha\coloneqq 24\sqrt{\frac{e(e+1)}{2}}$, and $\bar{\eta}=\frac{r}{Ld}$. Further define
@@
Let $\mathcal{F}_{k}:=\sigma(g_{0},P_{0},\ldots,g_{k-1},P_{k-1})$ and $\bar{\sigma}\coloneqq\alpha\cdot\frac{\sigma}{\sqrt{\bar{B}}}$. Recall that Proposition B.4 provides, for $|\lambda|\leq 1/\bar{\sigma}$,
@@
(a)  uses (40), (b)  uses $\|P_{k}^{\top}n_{k}\|\leq\|P_{k}^{\top}\|\|n_{k}\|=\sqrt{\frac{d}{r}}\|n_{k}\|$, and (c)  applies Proposition B.4 with the choice (41) (which is valid under (42)).
@@ Appendix C Missing Proofs for RS-NSGD
Proposition C.1 (Restatement of Proposition 5.1).
@@
Proposition C.2 (Restatement of Proposition 5.2).
@@
Suppose that $F$ satisfies Assumptions 2.1 and 2.2 , and that the gradient satisfies Assumptions 2.3 and 2.6. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-NSGD iteration. We set
@@
Since $F$ satisfies Assumption 2.2, we have
@@
Here, (a) follows from Proposition C.2 on the event $\{\|g_{k}\|\neq 0\}$ (using $\|g_{k}\|\neq 0\Rightarrow\|P_{k}^{\top}g_{k}\|\neq 0$ a.s., hence $G_{k}=\frac{P_{k}P_{k}^{\top}g_{k}}{\|P_{k}^{\top}g_{k}\|}$ a.s.), and from $G_{k}=0$ a.s. on $\{\|g_{k}\|=0\}$. For (b), note that on $\{\|g_{k}\|\neq 0\}$,
@@
Case 2: $\|g_{k}\|\neq 0$. Then $\|P_{k}^{\top}g_{k}\|\neq 0$ almost surely, so $G_{k}=\frac{P_{k}P_{k}^{\top}g_{k}}{\|P_{k}^{\top}g_{k}\|}$ almost surely, and by Proposition C.1,
@@
Following the same argument as in Hübler et al. [11], Assumption 2.6 implies
@@
Suppose that $F$ satisfies Assumptions 2.1 and 2.2. Then
@@
Suppose that $F$ satisfies Assumptions 2.1 and 2.2, and that the gradient satisfies Assumptions 2.3 and 2.6. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-NSGD iteration. We set
@@
Finally, using Lemma C.4, $\|\nabla F(x_{0})\|\leq\sqrt{2\Delta_{0}L}$, and $12\sqrt{2}\leq 17$, together with the definitions of $\bar{\eta}$ and $\bar{B}$, the stated bound follows. ∎
@@ Appendix D RS-NGD: Update Rule and Convergence Analysis
Assume that $F$ satisfies Assumptions 2.1 and 2.2. Let $\{x_{k}\}_{k\geq 0}$ be generated by RS-NGD from $x_{0}$. We set
@@
Next, the Haar property in Proposition C.1, in particular (61), implies that
@@
Moreover, by Proposition C.2 and the tower property,
@@
Assume that $F$ satisfies Assumptions 2.1 and 2.2. Let $\{x_{k}\}_{k\geq 0}$ be generated by RS-NGD from $x_{0}$. Fix $T\geq 1$ and set
@@
Finally, by Lemma C.4, we have

### v2

@@ 1 Introduction
Table 1: Convergence guarantees for stochastic gradient methods with randomized subspace updates (RS) and direction normalization (Norm). Exp./w.h.p.: in expectation/high probability. Noise: sub-Gaussian, BV, and $p$-BCM; see Sections 2.3, 2.3 and 2.3, respectively.
@@
However, these light-tailed noise conditions can be overly restrictive in modern machine-learning workloads, where stochastic gradients are often heavy-tailed [27]; this limitation is not caused by the randomized subspace projection but rather reflects the well-known sensitivity of SGD-type updates to heavy-tailed gradient noise [23, 10]. To overcome this issue, we leverage normalized stochastic gradient methods (NSGD), which admit both in-expectation and high-probability convergence guarantees under the weaker bounded $p$-th moment assumption stated later in Section 2.3 [11]. Unlike clipping-based methods, whose analyses typically require problem- or noise-dependent parameters, including the tail index $p$ [34, 22], NSGD can be analyzed without knowing $p$.
@@ 2.3 Problem Setting and Assumptions
While Section 2.3 is standard, it can be too restrictive in regimes where the stochastic gradient noise is heavy-tailed. To accommodate such behavior, we allow the noise to have only a bounded $p$-th moment for some $p\in(1,2]$.
@@
These noise assumptions are related as follows: [^3] Reason. Section 2.3 implies Section 2.3 by Jensen’s inequality. Section 2.3 is Section 2.3 with $p=2$, and Section 2.3 for some $p$ implies Section 2.3 for any $q\leq p$ (again by Jensen). Hence, Section 2.3 implies Section 2.3 for every $p\in(1,2]$.
@@ 4 High-Probability Convergence of RS-SGD
Assume that the stochastic gradient $\nabla f(x,\xi)$ satisfies Sections 2.3 and 2.3. For a fixed $x\in\mathbb{R}^{d}$, define the mini-batch gradient estimator
@@
By Section 4, increasing the mini-batch size reduces the noise, and the following theorem establishes convergence.
@@
Suppose that Sections 2.3, 2.3, 2.3 and 2.3 hold. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by RS-SGD. Let $\bar{B}\coloneqq\lceil\max\{1,BT^{q}\}\rceil$ for $B>0$ and $q>0$, $\alpha\coloneqq 24\sqrt{\frac{e(e+1)}{2}}$, and $\bar{\eta}=\frac{r}{Ld}$. Further define
@@ 5.1 In-expectation Bound
When bounding the first term in Section 5.1 by $\|S\|$, the relationship between $\operatorname{tr}(S)$ and $\|S\|$ becomes important. To this end, we introduce the following quantity.
@@
Suppose that Sections 2.3, 2.3, 2.3 and 2.3 hold. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by RS-NSGD. We set
@@ Appendix B Missing Proofs for RS-SGD
Assume $\mathbb{E}[X]=0$ and $\|X\|$ is $\sigma$-sub-Gaussian in the sense of Appendix B. Then for all $u\in\mathbb{R}^{d}$,
@@
By Appendix B with $\lambda=1/\sigma$, $\mathbb{E}[\exp(\|X\|^{2}/\sigma^{2})]\leq e$. Therefore,
@@
Proposition B.4 (Generalization of Section 4).
@@
By Appendix B, for each $i$ and all $u\in\mathbb{R}^{d}$,
@@
Thus, the assumption of Appendix B holds for $\bar{X}$ with
@@
Applying Appendix B yields
@@
Assume $F$ is $L$-smooth and bounded below by $F^{\ast}$. Assume $\mathbb{E}[\nabla f(x,\xi)]=\nabla F(x)$ and that $\|\nabla f(x,\xi)-\nabla F(x)\|$ is $\sigma$-sub-Gaussian (in the sense of Appendix B). Let $P_{k}$ be i.i.d. Haar random matrices, and let $\{x_{k}\}_{k\geq 0}$ be generated by the update rule
@@
Suppose that $F$ satisfies Sections 2.3 and 2.3, and that the gradient satisfies Sections 2.3 and 2.3. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-SGD iteration. Let $\bar{B}\coloneqq\lceil\max\{1,BT^{q}\}\rceil$ for $B>0$ and $q>0$, $\alpha\coloneqq 24\sqrt{\frac{e(e+1)}{2}}$, and $\bar{\eta}=\frac{r}{Ld}$. Further define
@@
Let $\mathcal{F}_{k}:=\sigma(g_{0},P_{0},\ldots,g_{k-1},P_{k-1})$ and $\bar{\sigma}\coloneqq\alpha\cdot\frac{\sigma}{\sqrt{\bar{B}}}$. Recall that Appendix B provides, for $|\lambda|\leq 1/\bar{\sigma}$,
@@
(a)  uses (40), (b)  uses $\|P_{k}^{\top}n_{k}\|\leq\|P_{k}^{\top}\|\|n_{k}\|=\sqrt{\frac{d}{r}}\|n_{k}\|$, and (c)  applies Appendix B with the choice (41) (which is valid under (42)).
@@ Appendix C Missing Proofs for RS-NSGD
Proposition C.1 (Restatement of Section 5).
@@
Proposition C.2 (Restatement of Section 5.1).
@@
Suppose that $F$ satisfies Sections 2.3 and 2.3 , and that the gradient satisfies Sections 2.3 and 2.3. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-NSGD iteration. We set
@@
Since $F$ satisfies Section 2.3, we have
@@
Here, (a) follows from Appendix C on the event $\{\|g_{k}\|\neq 0\}$ (using $\|g_{k}\|\neq 0\Rightarrow\|P_{k}^{\top}g_{k}\|\neq 0$ a.s., hence $G_{k}=\frac{P_{k}P_{k}^{\top}g_{k}}{\|P_{k}^{\top}g_{k}\|}$ a.s.), and from $G_{k}=0$ a.s. on $\{\|g_{k}\|=0\}$. For (b), note that on $\{\|g_{k}\|\neq 0\}$,
@@
Case 2: $\|g_{k}\|\neq 0$. Then $\|P_{k}^{\top}g_{k}\|\neq 0$ almost surely, so $G_{k}=\frac{P_{k}P_{k}^{\top}g_{k}}{\|P_{k}^{\top}g_{k}\|}$ almost surely, and by Appendix C,
@@
Following the same argument as in Hübler et al. [11], Section 2.3 implies
@@
Suppose that $F$ satisfies Sections 2.3 and 2.3. Then
@@
Suppose that $F$ satisfies Sections 2.3 and 2.3, and that the gradient satisfies Sections 2.3 and 2.3. Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by the RS-NSGD iteration. We set
@@
Finally, using Appendix C, $\|\nabla F(x_{0})\|\leq\sqrt{2\Delta_{0}L}$, and $12\sqrt{2}\leq 17$, together with the definitions of $\bar{\eta}$ and $\bar{B}$, the stated bound follows. ∎
@@ Appendix D RS-NGD: Update Rule and Convergence Analysis
Assume that $F$ satisfies Sections 2.3 and 2.3. Let $\{x_{k}\}_{k\geq 0}$ be generated by RS-NGD from $x_{0}$. We set
@@
Next, the Haar property in Appendix C, in particular (61), implies that
@@
Moreover, by Appendix C and the tower property,
@@
Assume that $F$ satisfies Sections 2.3 and 2.3. Let $\{x_{k}\}_{k\geq 0}$ be generated by RS-NGD from $x_{0}$. Fix $T\geq 1$ and set
@@
Finally, by Appendix C, we have